OpenStudy (anonymous):

what are the 2 solution sets of b^2-4b=0 .... I need help thnx

OpenStudy (anonymous):

b^2-4b=b(b-4)=0 then b=0 .........or..........b-4=0

OpenStudy (anonymous):

that was simple thnx

OpenStudy (anonymous):

welcome